OpenStudy (anonymous):

Abstract Algebra: The practice exam problem I have is "Prove that there are consecutive integers in a string of any desired (finite) length, all of which are not prime."

OpenStudy (anonymous):

I should note that this is in the same section as Euclids Algorithm so I think the proof may involve use a form of that, but I really have no idea how to go about doing this.

OpenStudy (anonymous):

My best idea so far is that this will involve a proof by contradiction, but, that I think will require a formula for generating an interval of length k between two primes, which seems like it will be impossible to just make up. Any thoughts from you clever math folks?

OpenStudy (zarkon):

I have thought of a way that involves \(n!\)

OpenStudy (zarkon):

as a short example look at 5! 5!+2=[5*4*3+1]*2 5!+3=[5*4*2+1]*3 5!+4=[5*3*2+1]*4 5!+5=[4*3*2+1]*5 here is 4 consecutive numbers that are not prime

OpenStudy (anonymous):

Hmmm, I think I follow. Is the idea then that you can chose any k for k! and add up to k?

OpenStudy (zarkon):

k!+2 k!+3 ... k!+k at least (k-1) consecutive non-primes

OpenStudy (anonymous):

Ok thats what I thought you mean. I was testing numbers and saw 5!+7 is 127, a prime, so I wanted to make sure that we defined a stopping point. I am amazed you just came up with that. I was reading the wikipedia on the prime gap trying to figure this out. Is the idea then a proof by induction for k+1? If so I can probably handle that.

OpenStudy (zarkon):

I don't think induction is needed....just choose any k...it clearly gives you at least k-1 consecutive nonprimes

OpenStudy (anonymous):

Hah, quite right. Thanks very much for taking a look!